What is the probability of dealing two red cards from a standard deck of cards?
Round to the nearest thousands
1
Expert's answer
2019-04-08T13:21:34-0400

A standard deck of cards has 52 cards. There are 26 red cards among them. The probability that the first and the second cards are red


"P={26 \over 52}*{{26-1} \over {52-1}}={25 \over 102}=0.245."

Or

Number of ways to choose 2 red cards from 26 red cards


"\binom{26}{2}={26! \over 2!(26-2)!}=325"

Number of ways to choose 2 cards from 52 cards


"\binom{52}{2}={52! \over 2!(52-2)!}=1326"

The probability of dealing two red cards from a standard deck of cards


"P=\binom{26}{2}/\binom{52}{2}={325 \over 1326}=0.245"